What is the magnitude of the electric field at a field point 4.0 m from a point charge 10.0 nC


1
Expert's answer
2020-12-17T07:28:51-0500

The magnitude of electric field is


"E=\frac{1}{4\pi\epsilon_0}\frac{q}{r^2}=5.6\text{ V/m}."